In Japan, 3.75% of the people on the average afflicted with COVID 19 die from it. In the Philippines 12
out of 150 recorded positive in the COVID 19 die from it, do you think that our government handle the
situation properly and effectively? Use 0.05 level of significance.
1
Expert's answer
2020-03-17T15:44:52-0400

To answer the question, a one proportion test should be used.

Null hypothesis "H_0: p=0.0375."

Alternative hypothesis "H_a: p>0.0375."

Test statistic: "z=\frac{\frac{12}{150}-0.0375}{\sqrt{\frac{0.0375(1-0.0375)}{150}}}=2.74."

P-value: "p=0.0031."

Since the P-value is less than 0.05, reject the null hypothesis.

There is a sufficient evidence that the government does not handle the

situation properly and effectively.